# Efficient quantum Gibbs samplers with Kubo–Martin–Schwinger detailed balance condition

### v2

@@ Appendix D Quadrature error analysis
Estimate the integral of $g$. Similarly, by Eq. D.6 with 15 and $|\tanh((\nu^{\prime}-\nu)/4)|\leq 1$, we have $\left\lVert g(t,t^{\prime})\right\rVert_{L^{\infty}(\mathbb{R}^{2})}\leq\frac{1}{2\pi^{2}}A_{q}^{4}S^{2}$. A direct computation gives, for a fixed $\nu^{\prime}\in\mathbb{R}$,
$\frac{\mathrm{d}}{\mathrm{d}\nu}\tanh((\nu^{\prime}-\nu)/4)=\frac{1}{4}(\tanh^{2}((\nu^{\prime}-\nu)/4)-1)\in L^{1}(\mathbb{R})\bigcap L^{\infty}(\mathbb{R})$
with its $L^{1}$ and $L^{\infty}$ norms independent of $\nu^{\prime}\in\mathbb{R}$. This implies
$\partial_{\nu^{\prime}}\left(\tanh((\nu^{\prime}-\nu)/4)\right)\partial_{\nu}\left(e^{-\nu/4}u(\nu)\right)\in L^{1}(\mathbb{R}^{2})\,.$
In addition, we have the second-order partial derivative:
$\partial_{\nu^{\prime}\nu}\tanh((\nu^{\prime}-\nu)/4)=\frac{1}{8}\tanh((\nu^{\prime}-\nu)/4)(1-\tanh^{2}((\nu^{\prime}-\nu)/4))\in L^{1}(\mathbb{R}^{2})\,,$
which is $L^{1}$-integrable in $\nu$ (for fixed $\nu^{\prime}$). Then, we can estimate
$\left\lVert t^{\prime}tg(t,t^{\prime})\right\rVert_{L^{\infty}(\mathbb{R}^{2})}$ $=\mathcal{O}\left(\iint_{\mathbb{R}^{2}}\left|\partial_{\nu^{\prime}\nu}\widehat{g}(\nu,\nu^{\prime})\right|\,\mathrm{d}\nu\,\mathrm{d}\nu^{\prime}\right)$
$=\mathcal{O}\left(A_{q}^{2}C_{1,u}^{2}+A_{q}^{4}S^{2}A_{w}^{2}+A_{q}^{3}C_{1,u}SA_{w}+A_{q}^{4}\beta S\right)\,.$
by a straightforward computation:
$\partial_{\nu^{\prime}\nu}\left(\tanh(\beta(\nu^{\prime}-\nu)/4)e^{-\beta(\nu+\nu^{\prime})/4}u(\beta\nu)w(\nu)\overline{u(\beta\nu^{\prime})w(\nu^{\prime})}\right)$
$=$ $\underbrace{\partial_{\nu^{\prime}\nu}\left(\tanh(\beta(\nu^{\prime}-\nu)/4)\right)\overline{w(\nu^{\prime})}}_{\left\lVert\cdot\right\rVert_{L^{1}}=\mathcal{O}(A_{q}S\beta)}\underbrace{e^{-\beta\nu/4}u(\beta\nu)w(\nu)\overline{u(\beta\nu^{\prime})e^{-\beta\nu^{\prime}/4}}}_{\left\lVert\cdot\right\rVert_{L^{\infty}}\leq A_{q}^{3}}$
$+$ $\underbrace{\partial_{\nu^{\prime}}\left(\tanh(\beta(\nu^{\prime}-\nu)/4)\right)\partial_{\nu}\left(e^{-\beta\nu/4}u(\beta\nu)\right)}_{\left\|\cdot\right\|_{L^{1}}=\mathcal{O}\left(C_{1,u}\right)}\underbrace{w(\nu)\overline{u(\beta\nu^{\prime})e^{-\beta\nu^{\prime}/4}w(\nu^{\prime})}}_{\|\cdot\|_{L^{\infty}}\leq A^{3}_{q}}$
$+$ $\underbrace{\partial_{\nu}\left(\tanh(\beta(\nu^{\prime}-\nu)/4)\right)\partial_{\nu^{\prime}}\left(\overline{u(\beta\nu^{\prime})e^{-\beta\nu^{\prime}/4}}\right)}_{\left\|\cdot\right\|_{L^{1}}=\mathcal{O}(C_{1,u})}\underbrace{e^{-\beta\nu/4}u(\beta\nu)w(\nu)\overline{w(\nu^{\prime})}}_{\left\|\cdot\right\|_{L^{\infty}}\leq A^{3}_{q}}$
$+$ $\underbrace{\left(\tanh(\beta(\nu^{\prime}-\nu)/4)\right)}_{\|\cdot\|_{L^{\infty}}=\mathcal{O}(1)}\underbrace{\partial_{\nu}\left(e^{-\beta\nu/4}u(\beta\nu)\right)\partial_{\nu^{\prime}}\left(\overline{u(\beta\nu^{\prime})e^{-\beta\nu^{\prime}/4}}\right)}_{\left\|\cdot\right\|_{L^{1}}\leq C^{2}_{1,u}}\underbrace{w(\nu)\overline{w(\nu^{\prime})}}_{\left\|\cdot\right\|_{L^{\infty}}\leq A^{2}_{q}}$
$+$ $\underbrace{\partial_{\nu}\left(\tanh(\beta(\nu^{\prime}-\nu)/4)e^{-\beta\nu/4}u(\beta\nu)\overline{u(\beta\nu^{\prime})e^{-\beta\nu^{\prime}/4}}\right)\partial_{\nu^{\prime}}\left(\overline{w(\nu^{\prime})}\right)}_{\left\|\cdot\right\|_{L^{1}}=\mathcal{O}(C_{1,u}A^{2}_{q}A_{\omega}S+A^{3}_{q}A_{\omega}S)}\underbrace{w(\nu)}_{\left\|\cdot\right\|_{L^{\infty}}\leq A_{q}}$
$+\underbrace{\partial_{\nu^{\prime}}\left(\tanh(\beta(\nu^{\prime}-\nu)/4)\overline{u(\beta\nu^{\prime})e^{-\beta\nu^{\prime}/4}}u(\beta\nu)e^{-\beta\nu/4}\right)\left(\partial_{\nu}w(\nu)\right)}_{\left\|\cdot\right\|_{L^{1}}=\mathcal{O}(C_{1,u}A^{2}_{q}A_{\omega}S+A^{3}_{q}A_{\omega}S)}\underbrace{\overline{w(\nu^{\prime})}}_{\left\|\cdot\right\|_{L^{\infty}}\leq A_{q}}$
$+\underbrace{\tanh(\beta(\nu^{\prime}-\nu)/4)\overline{u(\beta\nu^{\prime})e^{-\beta\nu^{\prime}/4}}u(\beta\nu)e^{-\beta\nu/4}}_{\left\|\cdot\right\|_{L^{\infty}}\leq A^{2}_{q}}\underbrace{(\partial_{\nu}w(\nu))\partial_{\nu^{\prime}}\left(\overline{w(\nu^{\prime})}\right)}_{\left\|\cdot\right\|_{L^{1}}\leq A^{2}_{q}A^{2}_{\omega}S^{2}}\,.$
In the same manner as Eq. D.10, we have
(D.15) $\int_{-T}^{T}\int_{-T}^{T}|g(t,t^{\prime})|\,\mathrm{d}t\,\mathrm{d}t^{\prime}=\mathcal{O}(A_{q}^{2}C_{1,u}^{2}+A_{q}^{4}S^{2}A_{w}^{2}+A_{q}^{3}C_{1,u}SA_{w}+A_{q}^{4}\beta S)\log^{2}(T)\,,\quad T>1\,.$
@@
(D.16) $T_{g}(A_{u},A_{w}):=\inf\left\{T>0\,;\ u^{2s-1}\leq\exp\left(\frac{s}{2(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}u\right)\ \text{for any $u\geq T^{1/s}$}\right\},$
which satisfy, as $2\beta A_{u}+2A_{w}+\beta\to\infty$,
$T_{g}(A_{u},A_{w})=\Theta\left((2\beta A_{u}+2A_{w}+\beta)\log\left(2\beta A_{u}+2A_{w}+\beta\right)\right)\,.$
It follows from Eqs. D.15 and D.2 and some similar estimates as in Eq. D.14 that
$\iint_{\mathbb{R}^{2}}|g(t,t^{\prime})|\,\mathrm{d}t\,\mathrm{d}t^{\prime}$ $=\int_{-T_{g}}^{T_{g}}\int_{-T_{g}}^{T_{g}}|g(t,t^{\prime})|\,\mathrm{d}t\,\mathrm{d}t^{\prime}+\mathcal{O}\left(A^{4}_{q}S^{2}\int_{T_{g}}^{\infty}\exp\left(-\frac{s}{(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}r^{1/s}\right)r\,\mathrm{d}r\right)$
$=\int_{-T_{g}}^{T_{g}}\int_{-T_{g}}^{T_{g}}|g(t,t^{\prime})|\,\mathrm{d}t\,\mathrm{d}t^{\prime}+\mathcal{O}\left(A^{4}_{q}S^{2}\int_{T_{g}^{1/s}}^{\infty}\exp\left(-\frac{s}{(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}u\right)u^{2s-1}\,\mathrm{d}u\right)$
$=\mathcal{O}\left((A_{q}^{2}C_{1,u}^{2}+A_{q}^{4}S^{2}A_{w}^{2}+A_{q}^{3}C_{1,u}SA_{w}+A_{q}^{4}\beta S)\log^{2}(\beta A_{u}+A_{w}+\beta)\right)\,.$
The proof is complete. ∎
We next prove 16. We shall employ the Poisson summation formula recalled in the following lemma [Pin08, Theorem 4.4.2].
@@
We are now ready to show 16.
Proof of 16.
In the proof, we shall consider an infinite time grid defined by $\{m\tau\}_{m\in\mathbb{Z}}$. The grid $\{t_{m}\}_{m=0}^{2M-1}$ is given by the subset $\{m\tau\}_{m=-M}^{M-1}$ as in the statement of 16. We omit the upper index $a$ in $f^{a}$ and $g^{a}$.
@@
$\leq$ $\sum_{|m|\geq M}\frac{A^{2}_{q}S}{\pi}\exp\left(\frac{es}{2}-\frac{s}{(\beta A_{u}+A_{w})^{1/s}e}\left|m\tau\right|^{1/s}\right)\tau\,.$
@@
(D.18) $\sum_{|m|\geq M}\exp\left(-\frac{s}{(\beta A_{u}+A_{w})^{1/s}e}\left|m\tau\right|^{1/s}\right)\tau$
$\leq$ $2\int^{\infty}_{(M-1)\tau}\exp\left(-\frac{s}{(\beta A_{u}+A_{w})^{1/s}e}\left|t\right|^{1/s}\right)\mathrm{d}t$
$\leq$ $2s\int^{\infty}_{((M-1)\tau)^{1/s}}u^{s-1}\exp\left(-\frac{s}{(\beta A_{u}+A_{w})^{1/s}e}u\right)\mathrm{d}u\,.$
Let $T_{f}(A_{u},A_{w})$ be the constant defined as in Eq. D.12. Then, when $(M-1)\tau\geq T_{f}(A_{u},A_{w})$, we can compute, by (D.18),
$\sum_{|m|\geq M}\exp\left(-\frac{s}{(\beta A_{u}+A_{w})^{1/s}e}\left|m\tau\right|^{1/s}\right)\tau$ $\leq 2s\int^{\infty}_{((M-1)\tau)^{1/s}}\exp\left(-\frac{s}{2(\beta A_{u}+A_{w})^{1/s}e}u\right)\mathrm{d}u$
$\leq 4s\frac{(\beta A_{u}+A_{w})^{1/s}e}{s}\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta A_{u}+A_{w})^{1/s}e}\right)\,.$
@@
$\leq$ $\sum_{|m|\geq M}\frac{A^{2}_{q}S}{\pi}\exp\left(\frac{es}{2}-\frac{s}{(\beta A_{u}+A_{w})^{1/s}e}\left|m\tau\right|^{1/s}\right)\tau$
$\leq$ $C_{s}A_{q}^{2}S(\beta A_{u}+A_{w})^{1/s}\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta A_{u}+A_{w})^{1/s}e}\right)\,.$
@@
$\leq$ $\left\|\int_{-\infty}^{\infty}f(t)A^{a}(t)\,\mathrm{d}t-\sum^{\infty}_{m=-\infty}f(m\tau)A^{a}(m\tau)\tau\right\|+C_{f}\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta A_{u}+A_{w})^{1/s}e}\right)\,,$
@@
$C_{f}=\mathcal{O}\left(A_{q}^{2}S(\beta A_{u}+A_{w})^{1/s}\right)\,.$
Note from (3.4) that $\int^{\infty}_{-\infty}f(t)A^{a}(t)\,\mathrm{d}t=\sum_{i,j}\widehat{f}(\lambda_{i}-\lambda_{j})P_{i}A^{a}P_{j}$, and from 15 that
@@
For $\tau<\frac{2\pi}{2\left\lVert H\right\rVert+S}$, by Poisson summation formula in 31, we have
@@
$\left\|L_{a}-\sum^{2M-1}_{m=0}f(t_{m})A^{a}(t_{m})\tau\right\|\leq C_{f}\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta A_{u}+A_{w})^{1/s}e}\right)\,.$
and concludes the proof of (3.15).
Estimate for the coherent term $G$. Again by 30 with similar estimates as in (D.19) for $L_{a}$, we obtain, when $(M-1)\tau>T_{g}(A_{u},A_{w})$,
$\left\|\sum^{\infty}_{n,m=-\infty}g(n\tau,m\tau)A^{a}(n\tau)A^{a}(m\tau)\tau^{2}-\sum^{2M-1}_{n,m=0}g(t_{n},t_{m})A^{a}(t_{n})A^{a}(t_{m})\tau^{2}\right\|$
$\leq$ $\|A^{a}\|^{2}\sum_{|n|,|m|\geq M}\left|g(n\tau,m\tau)\right|\tau^{2}\leq C_{g}\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}\right)\,,$
where the constant $T_{g}(A_{u},A_{w})$ is given in Eq. D.16 and
$C_{g}=\mathcal{O}\left(A_{q}^{4}S^{2}(2\beta A_{u}+2A_{w}+\beta)^{1/s}\right)\,.$
@@
(D.21) $\left\|\iint_{\mathbb{R}^{2}}g(t,t^{\prime})A^{a}(t)A^{a}(t^{\prime})\,\mathrm{d}t\,\mathrm{d}t^{\prime}-\sum^{2M-1}_{n,m=0}g(t_{n},t_{m})A^{a}(t_{n})A^{a}(t_{m})\tau^{2}\right\|$
$\leq$ $\left\|\iint_{\mathbb{R}^{2}}g(t,t^{\prime})A^{a}(t)A^{a}(t^{\prime})\,\mathrm{d}t\,\mathrm{d}t^{\prime}-\sum^{\infty}_{n,m=-\infty}g(n\tau,m\tau)A^{a}(n\tau)A^{a}(m\tau)\tau^{2}\right\|$
$+C_{g}\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}\right)\,.$
Similarly, Poisson summation formula with $\mathrm{supp}(\widehat{g})\subset[-S,S]\times[-S,S]$ gives
$\widehat{g}(\nu,\nu^{\prime})=\sum^{\infty}_{n,m=-\infty}\widehat{g}\left(\nu+\frac{2\pi n}{\tau},\nu^{\prime}+\frac{2\pi m}{\tau}\right)=\sum^{\infty}_{n,m=-\infty}g(n\tau,m\tau)e^{in\nu\tau}e^{-im\nu^{\prime}\tau}\tau^{2}\,,$
for $\tau<\frac{2\pi}{2\left\lVert H\right\rVert+S}$ and $(\nu,\nu^{\prime})\in B_{H}\times B_{H}$, thanks to $|\nu|,|\nu^{\prime}|\leq 2\|H\|$. It follows that the first term of (D.21) is zero. By definition (3.7) of $G$ and above estimates, it holds that
$\left\|G-\sum_{a\in\mathcal{A}}\sum^{2M-1}_{n,m=0}g(t_{n},t_{m})A^{a}(t_{n})A^{a}(t_{m})\tau^{2}\right\|$
$\leq$ $\sum_{a\in\mathcal{A}}\left\|\iint_{\mathbb{R}^{2}}g(t,t^{\prime})A^{a}(t)A^{a}(t^{\prime})\,\mathrm{d}t\,\mathrm{d}t^{\prime}-\sum^{2M-1}_{n,m=0}g(t_{n},t_{m})A^{a}(t_{n})A^{a}(t_{m})\tau^{2}\right\|$
$\leq$ $C_{g}|\mathcal{A}|\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}\right)\,.$
@@
Under 15, for any given $T>0$, there exists small $\tau=\Theta\left(1/A_{q}^{4}T^{2}S^{3}\right)$ such that for any integer $M$ with $M\tau\leq T$,
@@
$\left|f^{\prime}(t)\right|=\left|\frac{1}{2\pi}\int_{\mathbb{R}}u(\beta\nu)w(\nu)e^{-\beta\nu/4}\nu e^{-it\nu}\,\mathrm{d}\nu\right|\leq\frac{A_{q}}{2\pi}\int_{\mathbb{R}}|\nu w(\nu)|\,\mathrm{d}\nu=\mathcal{O}(A_{q}^{2}S^{2})\,.$
@@
$\left|\partial_{t}g(t,t^{\prime})\right|=\mathcal{O}(A_{q}^{4}S^{3})\,,\quad\left|\partial_{t^{\prime}}g(t,t^{\prime})\right|=\mathcal{O}(A_{q}^{4}S^{3})\,.$
@@
$\left|\sum^{M}_{m=-M}|f(m\tau)|\,\tau-\int^{(M+1)\tau}_{-M\tau}\left|f(t)\right|\,\mathrm{d}t\right|=\mathcal{O}\left(A_{q}^{2}S^{2}M\tau^{2}\right)\,,$

### v4

@@ Appendix D Error analysis and complexity
Estimate the integral of $g$. Similarly, by Eq. D.6 with 13 and $|\tanh(\nu/4)|\leq 1$, we have $\left\lVert g(t)\right\rVert_{L^{\infty}(\mathbb{R}^{2})}\leq\frac{2\xi_{q}S}{\pi}$. A direct computation gives
$\frac{\mathrm{d}}{\mathrm{d}\nu}\tanh(\nu/4)=\frac{1}{4}(\tanh^{2}(\nu/4)-1)\in L^{1}(\mathbb{R})\bigcap L^{\infty}(\mathbb{R})\,.$
Similar to (D.11), this gives
$2\pi|tg(t)|$ $\leq\int_{\mathbb{R}}\left|\widehat{g}^{\prime}(\nu)\right|\mathrm{d}\nu$
$\leq\frac{1}{2}\underbrace{\|\tanh(-\beta\nu/4)^{(1)}\|_{L^{1}(\mathbb{R})}}_{=\mathcal{O}(1)}\underbrace{\|w(\nu/4S)\|_{L^{\infty}(\mathbb{R})}}_{=\mathcal{O}(\xi_{q})}+\frac{1}{2}\underbrace{\|\tanh(-\beta\nu/4)\|_{L^{\infty}(\mathbb{R})}}_{=\mathcal{O}(1)}\underbrace{\|w(\nu/4S)^{(1)}\|_{L^{1}(\mathbb{R})}}_{=\mathcal{O}(1)}\,,$
namely,
$|g(t)|\leq\mathcal{O}(\xi_{q})/t\,.$
In the same manner as Eq. D.12, we have, for $T>\min\{1,1/S\}$
(D.17) $\int_{-T}^{T}|g(t)|\,\mathrm{d}t$ $=\int_{|t|\leq\min\{1,1/S\}}|g(t)|\,\mathrm{d}t+\int_{\min\{1,1/S\}\leq|t|\leq T}|g(t)|\,\mathrm{d}t$
$=\mathcal{O}(\xi_{q}(1+\log(T\max\{S,1\}))\,,$
@@
(D.18) $T_{g}:=\inf\left\{T>0\,;\ u^{s-1}\leq\exp\left(\frac{s}{2e(\beta+\xi_{w}/4S)^{1/s}}u\right)\ \text{for any $u\geq T^{1/s}$}\right\},$
which satisfy
$T_{g}=\Theta\left((\beta+\xi_{w}/4S)\log\left(\beta+\xi_{w}/4S\right)\right).$
Let $T=\Theta\left((\beta+\xi_{w}/4S)\log^{s}(\left(\beta+\xi_{w}/4S\right)\max\{S,1\})\right)>T_{g}$. It follows from Eqs. D.17 and D.2 and some similar estimates as in Appendix D that
$\int_{\mathbb{R}}|g(t)|\,\mathrm{d}t$ $=\int_{-T}^{T}|g(t)|\,\mathrm{d}t+\mathcal{O}\left(\int^{\infty}_{T^{1/s}}\xi_{q}S\exp\left(-\frac{s}{e\left(\beta+\xi_{w}/4S\right)^{1/s}}u\right)u^{s-1}\mathrm{d}u\right)$
$=\int_{-T}^{T}|g(t)|\,\mathrm{d}t+\mathcal{O}\left(\int^{\infty}_{T^{1/s}}\xi_{q}S\exp\left(-\frac{s}{2e\left(\beta+\xi_{w}/4S\right)^{1/s}}u\right)\mathrm{d}u\right)$
$=\mathcal{O}\left(\xi_{q}(1+\log((\beta+\xi_{w}/S)\max\{S,1\}))\right)\,.$
The proof is complete.
We next prove 15. We first restate it with more explicit parameter dependence.
Proposition 31 (restatement of 15).
Under 13, we assume $\beta>0$, $\|A^{a}\|\leq 1$ for any $a\in\mathcal{A}$. When $\tau\leq\frac{\pi}{\left\lVert H\right\rVert+2S}$ and
$(M-1)\tau=\Omega\left((\beta\xi_{u}+\xi_{w}/S)\log\left(\beta\xi_{u}+\xi_{w}/S\right)\right)\,,$
it holds that
(D.19) $\left\|L_{a}-\sum^{2M-1}_{m=0}f^{a}(t_{m})A^{a}(t_{m})\tau\right\|\leq C_{f}S\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}\right)\,,$
with
$C_{f}=\mathcal{O}\left(\xi_{q}^{2}(\beta\xi_{u}+\xi_{w}/S)^{1/s}\right)\,,$
and
$\left\|G-\sum^{2M-1}_{m=0}g(t_{m})H_{L}(t_{m})\tau\right\|\leq C_{g}S|\mathcal{A}|\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2e(\beta+\xi_{w}/4S)^{1/s}}\right)\,,$
with
$C_{g}={\mathcal{O}}\left(\xi_{q}(\xi_{q}C_{1,u}+\xi^{2}_{q}\xi_{w})^{2}\left(\beta+\xi_{w}/4S\right)^{1/s}\log^{2}((\beta\xi_{u}+\xi_{w}/S)\max\{S,1\})\right)\,.$
We shall employ the Poisson summation formula in the following lemma [Pin08, Theorem 4.4.2].
@@
We are now ready to show 31.
Proof of 31.
In the proof, we shall consider an infinite time grid defined by $\{m\tau\}_{m\in\mathbb{Z}}$. The grid $\{t_{m}\}_{m=0}^{2M-1}$ is given by the subset $\{m\tau\}_{m=-M}^{M-1}$ as in the statement of 31. We omit the upper index $a$ in $f^{a}$.
@@
$\leq$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\xi^{2}_{q}S}{\pi}\exp\left(\frac{es}{2}-\frac{s}{e}\left(\frac{|m\tau|}{\beta\xi_{u}+\xi_{w}/S}\right)^{1/s}\right)\tau}\,.$
@@
(D.21) $\sum_{|m|\geq M}\exp\left(-\frac{s}{(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}\left|m\tau\right|^{1/s}\right)\tau$
$\leq$ $2\int^{\infty}_{(M-1)\tau}\exp\left(-\frac{s}{(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}\left|t\right|^{1/s}\right)\mathrm{d}t$
$\leq$ $2s\int^{\infty}_{((M-1)\tau)^{1/s}}u^{s-1}\exp\left(-\frac{s}{(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}u\right)\mathrm{d}u\,.$
Let $T_{f}$ be defined as in Eq. D.14. Then, when $(M-1)\tau\geq T_{f}$, we can compute, by (D.21),
$\sum_{|m|\geq M}\exp\left(-\frac{s}{(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}\left|m\tau\right|^{1/s}\right)\tau$ $\leq 2s\int^{\infty}_{((M-1)\tau)^{1/s}}\exp\left(-\frac{s}{2(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}u\right)\mathrm{d}u$
$\leq 4s\frac{(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}{s}\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}\right)\,.$
@@
$\leq$ $\sum_{|m|\geq M}\frac{\xi^{2}_{q}S}{\pi}\exp\left(\frac{es}{2}-\frac{s}{(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}\left|m\tau\right|^{1/s}\right)\tau$
$\leq$ $C_{s}\xi_{q}^{2}S(\beta\xi_{u}+\xi_{w}/S)^{1/s}\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}\right)\,.$
@@
$\leq$ $\left\|\int_{-\infty}^{\infty}f(t)A^{a}(t)\,\mathrm{d}t-\sum^{\infty}_{m=-\infty}f(m\tau)A^{a}(m\tau)\tau\right\|+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{f}S\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}\right)}\,,$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{f}=\mathcal{O}\left(\xi_{q}^{2}(\beta\xi_{u}+\xi_{w}/S)^{1/s}\right)\,.}$
Note from (3.3) that $\int^{\infty}_{-\infty}f(t)A^{a}(t)\,\mathrm{d}t=\sum_{i,j}\widehat{f}(\lambda_{i}-\lambda_{j})P_{i}A^{a}P_{j}$, and from 13 that
@@
For $\tau<\frac{2\pi}{2\left\lVert H\right\rVert+S}$, by Poisson summation formula in 32, we have
@@
$\left\|L_{a}-\sum^{2M-1}_{m=0}f(t_{m})A^{a}(t_{m})\tau\right\|\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{f}S\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}\right)}\,.$
and concludes the proof of (D.19).
Estimate for the coherent term $G$. By the inequality (D.3) in 30, we obtain
$\|L_{a}\|\leq\int_{\mathbb{R}}|f(t)|\|A^{a}\|\,\mathrm{d}t=\mathcal{O}\left((\xi_{q}C_{1,u}+\xi^{2}_{q}\xi_{w})\log((\beta\xi_{u}+\xi_{w}/S)\max\{S,1\})\right)\,,$
which implies, recalling Eq. 3.11,
$\sup_{t\in\mathbb{R}}\|H_{L}(t)\|\leq\sum_{a\in\mathcal{A}}\|L_{a}\|^{2}=\mathcal{O}\left(|\mathcal{A}|(\xi_{q}C_{1,u}+\xi^{2}_{q}\xi_{w})^{2}\log^{2}((\beta\xi_{u}+\xi_{w}/S)\max\{S,1\})\right)\,.$
Again by 30 with similar estimates as in (D.22) for $L_{a}$, when $(M-1)\tau\geq T_{g}$ with $T_{g}$ defined in (D.18), there holds
$\left\|\sum^{\infty}_{m=-\infty}g(m\tau)H_{L}(m\tau)\tau-\sum^{2M-1}_{m=0}g(t_{m})H_{L}(t_{m})\tau\right\|$
$\leq$ $\sup_{t\in\mathbb{R}}\|H_{L}(t)\|\sum_{|m|\geq M}\left|g(m\tau)\right|\tau\leq C_{g}S|\mathcal{A}|\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2e(\beta+\xi_{w}/4S)^{1/s}}\right)\,,$
where
$C_{g}={\mathcal{O}}\left(\xi_{q}(\xi_{q}C_{1,u}+\xi^{2}_{q}\xi_{w})^{2}\left(\beta+\xi_{w}/4S\right)^{1/s}\log^{2}((\beta\xi_{u}+\xi_{w}/S)\max\{S,1\})\right)\,.$
@@
(D.24) $\left\|G-\sum^{2M-1}_{m=0}g(t_{m})H_{L}(t_{m})\tau\right\|$
$\leq$ $\left\|\int_{-\infty}^{\infty}g(t)H_{L}(t)\,\mathrm{d}t-\sum^{\infty}_{m=-\infty}g(t_{m})H_{L}(t_{m})\tau\right\|+\left\|\sum^{\infty}_{m=-\infty}g(t_{m})H_{L}(t_{m})\tau-\sum^{2M-1}_{m=0}g(t_{m})H_{L}(t_{m})\tau\right\|$
$\leq$ $\left\|\int_{-\infty}^{\infty}g(t)H_{L}(t)\,\mathrm{d}t-\sum^{\infty}_{m=-\infty}g(t_{m})H_{L}(t_{m})\tau\right\|+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{g}S|\mathcal{A}|\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2e(\beta+\xi_{w}/4S)^{1/s}}\right)}\,.$
For $\tau<\frac{2\pi}{2\left\lVert H\right\rVert+4S}$, again by Poisson summation formula in 32, the first term of (D.24) is zero. By definition (3.10) of $G$ and above estimates, it holds that
$\left\|G-\sum^{2M-1}_{m=0}g(t_{m})H_{L}(t_{m})\tau\right\|\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{g}S|\mathcal{A}|\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2e(\beta+\xi_{w}/4S)^{1/s}}\right)}\,.$
@@
Under 13, for any given $T>0$, there exists small $\tau=\Theta\left(1/\xi_{q}^{2}S^{2}T\right)$ such that for any integer $M$ with $M\tau\leq T$,
@@
$\left|f^{\prime}(t)\right|=\left|\frac{1}{2\pi}\int_{\mathbb{R}}u(\beta\nu)w(\nu/S)e^{-\beta\nu/4}\nu e^{-it\nu}\,\mathrm{d}\nu\right|\leq\frac{\xi_{q}}{2\pi}\int_{\mathbb{R}}|\nu w(\nu/S)|\,\mathrm{d}\nu=\mathcal{O}(\xi_{q}^{2}S^{2})\,.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left|g^{\prime}(t)\right|=\mathcal{O}(\xi_{q}S^{2})}\,.$
@@
$\left|\sum^{M}_{m=-M}|f(m\tau)|\,\tau-\int^{(M+1)\tau}_{-M\tau}\left|f(t)\right|\,\mathrm{d}t\right|=\mathcal{O}\left(\xi_{q}^{2}S^{2}M\tau^{2}\right)\,,$